OpenStudy (anonymous):

The area of a rectangle is given by the expression 6x^2-13x-5 a) What is the length of the rectangle if its width is (2x+5)? b) if x=9cm, find the length, width, and area of a rectangle PLS HELP ME!!! SHOW YOUR SOLUTION write also A or B its Factoring Special Products Pls help me

OpenStudy (anonymous):

You hate math Help!!

OpenStudy (anonymous):

I hate Miley Cyrus !! :P

OpenStudy (anonymous):

pls help me...

OpenStudy (anonymous):

but you hate math help lol I will just give me sometime

OpenStudy (anonymous):

pls just help

OpenStudy (anonymous):

Okay \[A = 6x^2 -13x-5\] Well I think your first option is insufficient so I will try b

OpenStudy (anonymous):

ishaan the maths superhero......;)

OpenStudy (anonymous):

When x=9 \[A = 6 \times 81 - 13 \times 9 -5\]

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

Now Solve it through Calculator

OpenStudy (akshay_budhkar):

6x^2-13x-5 divided by 2x+5 wil give you the width

OpenStudy (anonymous):

ummm thats b? the answer? ok

OpenStudy (anonymous):

OK Akshay is right !!!

OpenStudy (anonymous):

a is not wrong ishaan...................

OpenStudy (anonymous):

A=6x^2−13x−5 When x=9 A=6×81−13×9−5

OpenStudy (anonymous):

yeah Hash I didn't get that at first

OpenStudy (anonymous):

divide area by width to get length expression

OpenStudy (akshay_budhkar):

as area is length*width

OpenStudy (anonymous):

6x^2-13x-5 divided by 2x+5 wil give you the width pls solve for me thanks

OpenStudy (anonymous):

Yeah Hash and Akshay are right thers is no need for superhero :D

OpenStudy (akshay_budhkar):

why am i getting its not divisible??????

OpenStudy (anonymous):

pls solve this 6x^2-13x-5 divided by 2x+5 wil give you the width

OpenStudy (akshay_budhkar):

Hash786 ishaan??????? its not divisible u get a remainder

OpenStudy (anonymous):

it is a wrong quesstion.................there is a remainder

OpenStudy (anonymous):

which must not come particularly if A=LxB

OpenStudy (akshay_budhkar):

i believe its wrong ihatemathhelp

OpenStudy (anonymous):

what is the length of the rectangle if its width is (2x+5)?

OpenStudy (anonymous):

can u check the area expression i hate math

OpenStudy (akshay_budhkar):

still the same!! wrong ques..

OpenStudy (anonymous):

i will scan the retricecan u still help me? 1min? i will scan

OpenStudy (akshay_budhkar):

sure

OpenStudy (anonymous):

ok we can try to........

OpenStudy (anonymous):

w8

OpenStudy (anonymous):

Oh Lovely Akshay will do it and Hash too

OpenStudy (anonymous):

lol

OpenStudy (akshay_budhkar):

ishaan is not interested he wants tough ques right?

OpenStudy (anonymous):

w8 its scanning

OpenStudy (anonymous):

Well the thing is i hate calculation ..and theres a lot in it

OpenStudy (anonymous):

Just get x.....

OpenStudy (anonymous):

help me :(

OpenStudy (anonymous):

what do you mean? can u help ,me can u solve it for me? pls...

OpenStudy (akshay_budhkar):

estudier the function is not divisible.....:(

OpenStudy (anonymous):

I guess u mean not factorisable, it is of course, divisible, just with a remainder. However, considering the level of the question, I agree it is very likely a mistake in the question.